What is the factored form of the quadratic equation "x^2+x-12=0".I also was wondering if you could show me how to work it out or understand how to solve this problem.
Thanks,
Jenae

Answer by Nate(3500)

First, let me show you the standard form: ax^2 + bx + c = 0
The first step is to multiply 'a' and 'c'.
Next, find factors of that product that add up to give you 'b'.
Then, plug in those numbers and simplify with grouping and such.
Last, you write the answer down!
...........................................................................
x^2+x-12=0 original
x^2+x-12=0 the product of 'a' and 'c' is -12 and the factors that add up to make 'b' are 4 and -3
x^2+4x-3x-12=0 plug in the numbers
(x^2+4x)+(-3x-12)=0 group the set
x(x+4)-3(x+4)=0 reduce by taking out the common factor
(x-3)(x+4)=0 simplify with grouping
Now, use the zero product property:
x-3=0 is x=3
x+4=0 is x=-4